Tank A contains a mixture of 10 gallons water& 5 gallons pure alcohol.Tank B has 12 gallons water& 3 gallons pure alcohol.How many gallons should be taken from each tank and combined in order to obtain 8 gallon soln containing 25%alchl by vol...?
Answer by stanbon(75887)

Tank A contains a mixture of 10 gallons water & 5 gallons pure alcohol.
50% alcohol
------------------
Tank B has 12 gallons water& 3 gallons pure alcohol.
25% alcohol
------------------
How many gallons should be taken from each tank and combined in order to obtain 8 gallon solution containing 25% alcohol by vol...?
-------------------
Equation:
alcohol + alcohol = alcohol
0.50x + 0.25(8-x) = 0.25*8
----
You could work the equation and find x = 0
or you could see it's obvious that you
would take none of the 50% solution and 8 gallons
of the 25% solution.
---------------------------------
Here's the math:
0.50x + 0.25(8-x) = 0.25*8
Multiply thru by 100 to get:
50x + 25*8 - 25x = 25*8
25x = 0
x = 0 (amt. of 50% solution needed)
---
8-x = 8 gallons (amt. of 25% solution needed)
===================================================
Cheers,
Stan H.
============